narutouzamaki2:

What is the sum of the angle measures in a 16-gon? 1,800° 154.3° 150° 144

narutouzamaki2:

i dont understand

justus:

Here its asking you to find the sum of interior angles of a polygon. To find the sum, we use this formula: Sum = (a-2) x 180° in this case a is 16 given from the question thus, sum = (16-2) × 180° = 14 × 180° = ________